# Aperiodic points for outer billiards

### v3

@@ 1.2. Methods
For this problem — and here comes the principal novelty of this paper — we employ scissors congruence invariants. Two polygons are scissors congruent if one can cut one polygon into finitely many polygonal pieces and assemble the other polygon from these pieces. If $\mathsf{G}$ is a subgroup of $\mathrm{Isom}(\mathbb{R}^{2})$, the full isometry group of $\mathbb{R}^{2}$, whose elements are the only ones allowed to act on the pieces, then we speak of $\mathsf{G}$-scissors congruence. A case of special importance for us is $\mathsf{G}=\mathsf{T}$: the group of all translations. By ($\mathsf{G}$-)scissors congruence invariants, we mean functions on polygons that take the same values on $\mathsf{G}$-scissors congruent polygons. Dynamical versions of $\mathsf{T}$-scissors congruence invariants are conjugacy invariants of polygon exchange transformations. [^1] Polygon exchange transformations are themselves scissors congruences, and dynamical invariants are, in a sense, scissors congruence invariants of these scissors congruences.
Such invariants, in principle, can lead to very general necessary conditions of periodicity, in the form of vanishing of these invariants. A sample condition of this type follows from Theorem 2.29: for a periodic polygon exchange transformation, any dynamical valuation (as described in Section 2.5) must vanish. The principle idea of Theorem 1.1 is simple: compute suitable dynamical invariants and show they are non-zero. Once it is established that a bounded restriction of $f_{\Pi}$ is not periodic, establishing the existence of periodic points is not immediate, but this part is less conceptual and more technical.
Using earlier results (cf. [Ruk22]), the Main Theorem can be reduced to the case of even $N$: for an odd $N$, there is a certain correspondence between outer billiards on the $N$-gon and on the $2N$-gon, allowing to conclude that the existence of aperiodic points in one of these systems implies the same for the other. Thus, it suffices to assume that $N$ is even. We also need to reduce the dynamics of $f_{\Pi}$ to that of a polygon exchange transformation, i.e., a piecewise isometry acting by translations. One obvious choice is the second iterate $f^{2}_{\Pi}$, however, a different choice allows to significantly simplify the computations. Namely, consider the half-turn $M$ about the center of $\Pi$, and form the composition $g:=M\circ f_{\Pi}$. It acts as a piecewise translation, see Fig. 3. Also, an $f_{\Pi}$-invariant bounded domain is invariant under $M$ and hence under $g$, and the dynamics of $g$ is not much different from that of $f_{\Pi}$, since $f_{\Pi}$ commutes with $M$. For these reasons, if $g$ is not periodic, then neither is a bounded restriction of $f_{\Pi}$. To establish non-periodicity of $g$, we use a two-dimensional dynamical invariant based on the Hadwiger–Glur valuation [HG51]. Explicit computation of this invariant and verification that the result is nonzero are not hard.
@@
Even though the two approaches are, by the end of the day, rely on similar ideas (non-vanishing of dynamical valuations), their implementations look different, and the invariants employed are also different. Any one of the two approaches is enough to prove the main theorem, however, we describe both. This is done on purpose: one or the other approch may work better for other problems.
@@ 1.3. Organization of the paper
Section 2 discusses general concepts associated with scissors congruences, including their dynamical aspects. As a special case, we consider polygon exchange transformations, which act as piecewise translations. Dynamical valuations of polytope exchange transformations of a given polygon are certain homomorphisms of the group of such transformations to the additive group of a real vector space. It follows that these dynamical valuations are conjugacy invariants under orientation preserving piecewise Euclidean conjugacies, and, for this reason, they are also called dynamic invariants. Dynamic Hadwiger valuations can be obtained as special cases of more general dynamic invariants; they are based on (translation) scissors congruence invariants of Hadwiger and Glur [HG51] initially discovered as a means of solving translation scissors congruence problem for planar polygons.
Assuming that $N$ is even, Section 3 deals with the outer billiard map $f_{\Pi}$ on a regular $N$-gon $\Pi$. There is a bounded domain $X$ invariant under $f_{\Pi}$; see Section 3.1. As explained above, the map $g_{X}:=M\circ f_{\Pi}|_{X}$ is a polygon exchange transformation of a bounded domain. Here $M$ is the half-turn about the center of $\Pi$. Computing the dynamical Hadwiger invariant of $g_{X}$ (Theorem 3.6) and establishing that it is nonzero (Corollary 3.7) yields that $f_{\Pi}$ is not periodic on $X$ (Theorem 3.8). We still need to upgrade this result to existence of aperiodic points, which is done in Section 6.
@@ 1.4. Acknowledgements
Valery Zgursky was a member of the team at an earlier stage of the project and co-authored a previous version of the paper; with gratitude, we acknowledge his contribution to the overall understanding of the problem and intuition through a series of computer experiments. We are grateful to Victor Kleptsyn for useful discussions and for communicating a proof of Proposition 2.33.
@@ 2.1. Basic terminology
Given a scissors isomorphism $f$ of $\Delta$, define a boundary point of $f$ as a point $z\in\Delta$, for which there is $n\in\mathbb{Z}$ such that $z\notin\mathrm{dom}(f^{n})$, that is, some positive or negative iterate of $f$ is not defined at $z$ (which includes the case $z\in\partial\Delta$). Clearly, boundary points of $f$ are precisely boundary points of all forward or backward pieces of $f$. If $f^{n}(z)=z$ for some $n\neq 0$ (which means, in particular, that $z\in\mathrm{dom}(f^{n})$), then the point $z$ is called periodic under $f$. The smallest positive integer $n$ with this property is called the period of $z$. It is immediate from the definition that periodic points are not boundary points. Aperiodic points are defined as points of $\Delta$ that are neither periodic points nor boundary points.
@@ 2.2. Fatou and Julia sets
Components of the Fatou set may be either locally periodic (all non-boundary points are periodic) or wandering (an open domain $\Omega\subset F(f)$ is said to be wandering if $f^{n}(\Omega)\cap\Omega=\varnothing$ unless $n=0$). The latter case is of course impossible if $\Delta$ has finite area.
@@
Clearly, if $f$ is periodic, then there are no essential open sets, hence $J$ is empty. On the other hand, if $f$ is non-periodic, then there is either an aperiodic point, which necessarily lies in $J$, or a sequence of periodic points whose periods tend to infinity, and any limit point of this sequence lies in $J$. ∎
@@
Recall that $f$ preserved the Euclidean area, in particular, the Poincaré recurrence theorem is applicable to $f$. It says that all points of $\Delta$ are recurrent, except perhaps for a set of zero area.
@@
Suppose that $f$ is periodic of period $n_{*}$, then $\Gamma_{+}=\Gamma_{n_{*}}$. Conversely, assume that $\Gamma_{+}=\Gamma_{n_{*}}$ for some $n_{*}>0$. Then the complement of $\Gamma_{+}$ is forward invariant under $f$, hence all complementary components are periodic. It follows that the complement of $\Gamma_{+}$ is also backward invariant under $f$, that is, $\Gamma_{-}=\Gamma_{+}=\Gamma$, as claimed. ∎
@@ 2.3. Polygon exchange transformations
Let $\mathsf{T}$ be the group of all translations of the plane; $\mathsf{T}$-scissors automorphisms are also known as polygon exchange transformations. Note also that $\mathsf{T}$-scissors congruences are synonymous with translation scissors congruences. The notion of a polygon exchange transformation admits a straightforward generalization to higher dimensions; the corresponding general notion is called a polytope exchange transformation. Dimension one polytope exchange transformations are called interval exchange transformations. An interval exchange transformation can be defined on a straight line interval or a finite union of intervals.
Given a bounded polyhedron $\Delta$ of dimension $d$, we write $\mathrm{PET}(\Delta)$ for the group of polytope exchange transformations of $\Delta$, that is, the group of all automorphisms of $\Delta$ in the $\mathsf{T}$-scissors congruence groupoid. More precisely, elements of $\mathrm{PET}(\Delta)$ are classes of polytope exchange transformations with respect to the equivalence relation identifying two maps if they coincide except on a finite union of affine hyperplanes. Periodic polytope exchange transformations yield, in particular, periodic elements of the group $\mathrm{PET}(\Delta)$. If $d=1$, then one writes $\mathrm{IET}(\Delta)$ for the group of all interval exchange transformations of $\Delta$. Abbreviations IET and PET stand, of course, for “interval exchange transformations” and “polytope exchange transformations”, respectively. We concentrate on dimensions two or one in this paper even though some notions and results make sense in arbitrary dimensions.
@@
Let $I$ be an interval and $I_{1}$, $I_{2}\subset I$ be disjoint (except possibly a shared endpoint) subintervals of equal length. Then there is an interval exchange transformation $s$ that takes $I_{1}$ onto $I_{2}$, takes $I_{2}$ onto $I_{1}$, and leaves all other points of $I$ fixed. Call $s$ an interval swap map.
@@
A point of $\Delta$ can lie in the impressions of only finitely many nests, in fact, no more than $m_{*}$ nests, where $m_{*}$ is from Lemma 2.13.
@@
Lemma 2.13 also motivates the notion of a basic sector.
@@
    any other sector with property (1) coincides with $\Lambda_{z}$ or is disjoint from $\Lambda_{z}$.
@@
Here, $m_{*}$ is again the same as in Lemma 2.13. For example, if $z$ is not a boundary point, then it lies in the impression of only one nest (namely, the level $n$ component of this nest is recovered as the level $n$ bi-piece containing $z$), and the only elementary sector at $z$ is the entire plane.
We proceed by reviewing an invariant of an interval exchange transformation called the Sah–Arnoux–Fathi invariant (abbreviated as SAF invariant) [Arn81, Sah81]. Consider an interval exchange transformation $g$ defined on an interval $I$. Then the SAF invariant $\mathrm{Inv}(g)$ is an element of $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ defined as follows. If subintervals $I_{1}$, $\dots$, $I_{n}$ of lengths $\lambda_{1}$, $\dots$, $\lambda_{n}$, respectively, form a partition of $I$, and the restriction of $g$ to each $I_{i}$ is a translation by $t_{i}\in\mathbb{R}$, then
$\mathrm{Inv}(g)=\lambda_{1}\otimes t_{1}+\lambda_{2}\otimes t_{2}+\dots+\lambda_{n}\otimes t_{n}.$
@@
$\mathrm{Inv}_{\beta}(g):=\beta(\lambda_{1},t_{1})+\beta(\lambda_{2},t_{2})+\dots+\beta(\lambda_{n},t_{n}).$
@@
Next, we want to generalize the SAF invariant to higher dimensional polytope exchange transformations. Specifically, define a dynamical valuation as a group homomorphism from $\mathrm{PET}(P)$ to $\mathbb{R}$. In what follows, new dynamical valuations will be introduced in 2D (they also make sense in dimensions $>2$).
@@ 2.4. The ideal Julia set
A nest $(Q_{n})$ is said to be essential if $Q_{n}$ is an essential set, for every $n\geqslant 0$. Say that $(Q_{n})$ is periodic if $Q_{n+1}=Q_{n}$ for all sufficiently large $n$; otherwise $(Q_{n})$ is said to be non-periodic.
@@
All results about the ideal Julia set can be rephrased in terms of symbolic dynamics. Consider the alphabet $\mathcal{A}$ whose symbols are in one-to-one correspondence with components of $\mathrm{dom}(f)$; one can formally set $\mathcal{A}:=\pi_{0}(\mathrm{dom}(f))$. Every nest $(Q_{n})$ gives rise to a bi-infinite sequence in the alphabet $\mathcal{A}$, namely, the $n$-th symbol is the component of $\mathrm{dom}(f)$ containing $f^{n-1}(Q_{n})$ if $n>0$ and the component of $\mathrm{dom}(f)$ containing $f^{n}(Q_{-n+1})$ if $n\leqslant 0$. This correspondence embeds the dynamical system $(\mathcal{J},h_{f})$ into $(\mathcal{A}^{\mathbb{Z}},\sigma)$, where $\sigma$ is the shift map.
@@ 2.5. Valuations
A valuation on convex polygons with values in $\mathbb{R}$ is by definition a function $\mu:\mathcal{P}\to\mathbb{R}$ such that $\mu(\varnothing)=0$ and
@@
Fix an area form on $\mathbb{R}^{2}$. Then one can assign the area $\mathrm{area}(P)$ to any polytope $P$. Area is a simple valuation invariant under the full group of isometries of $\mathbb{R}^{d}$. Now let $\xi:\mathbb{R}\to\mathbb{R}$ be any $\mathbb{Q}$-linear function. Then $\xi\circ\mathrm{area}$ is also a simple isometry invariant valuation; this construction gives uncoutably many such valuations, linearly independent over $\mathbb{R}$. Perimeter (suitably extended to degenerate polygons) is also an isometry invariant valuation, however, it is not simple.
@@
Let $P$ be a bounded polygon. By a dynamical valuation, we mean a group homomorphism $\Psi:\mathrm{PET}(P)\to\mathbb{R}$, where $\mathbb{R}$ is viewed as a group under addition. For any two transformations $f$, $g\in\mathrm{PET}(P)$, we therefore have
@@
It follows that $\Psi(f^{n})=n\Psi(f)$ for every $n\in\mathbb{Z}$. Therefore, if $f$ is periodic, then any dynamical valuation must vanish on $f$. Sometimes, dynamical valuations are also referred to as dynamical invariants, since they are, indeed, congugacy invariants ($\Psi(g\circ f\circ g^{-1})=\Psi(f)$). Below, we give a general construction of multidimensional dynamical invariants.
@@
$\mathrm{Inv}_{\mu}(f)=\sum_{i=1}^{n}\mu(P_{i})\otimes\vec{v}_{i}\in\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{d}$
@@
In particular, any $\mathbb{Q}$-bilinear functional $\beta:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{C}$ defines a dynamical valuation $\mathrm{Inv}_{\mu,\beta}$ with values in $\mathbb{C}$, which can be computed as the sum of $\beta(\mu(P_{i}),\vec{v}_{i})$ for $i=1$, $\dots$, $n$.
@@
The dynamic Hadwiger invariant lies in the kernel of the multiplication map $\mathfrak{m}:\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{d}\to\mathbb{R}^{d}$ given by $\mathfrak{m}(x\otimes\vec{v})=x\vec{v}$.
@@
Let $f$ be a polytope exchange transformation of a bounded polygon $P$ so that $\mathrm{dom}(f)$ consists of open polygons $P_{1}$, $\dots$, $P_{n}$, and $f|_{P_{i}}$ is the translation by vector $\vec{v}_{i}$. Observe that $\mathfrak{M}(f):=\mathfrak{m}(\mathrm{Inv}_{L}(f))$ can be computed as the sum of $\mu_{L}(P_{i})\vec{v}_{i}$. We want to show that $\mathfrak{M}(f)=0$.

### v4

@@ 1.2. Methods
For this problem — and here comes the principal novelty of this paper — we employ scissors congruence invariants. Two polygons are scissors congruent if one can cut one polygon into finitely many polygonal pieces and assemble the other polygon from these pieces. If $\mathsf{G}$ is a subgroup of $\mathrm{Isom}(\mathbb{R}^{2})$, the full isometry group of $\mathbb{R}^{2}$, whose elements are the only ones allowed to act on the pieces, then we speak of $\mathsf{G}$-scissors congruence. A case of special importance for us is $\mathsf{G}=\mathsf{T}$: the group of all translations. By ($\mathsf{G}$-)scissors congruence invariants, we mean functions on polygons that take the same values on $\mathsf{G}$-scissors congruent polygons. Dynamical versions of $\mathsf{T}$-scissors congruence invariants are conjugacy invariants of polygon exchange transformations. [^1] Polygon exchange transformations are themselves scissors congruences, and dynamic invariants are, in a sense, scissors congruence invariants of these scissors congruences.
Such invariants, in principle, can lead to very general necessary conditions of periodicity, in the form of vanishing of these invariants. A sample condition of this type follows from Theorem 2.29: for a periodic polygon exchange transformation, any dynamic invariant (as described in Section 2.6) must vanish. The principle idea of Theorem 1.1 is simple: compute suitable dynamic invariants and show they are non-zero. Once it is established that a bounded restriction of $f_{\Pi}$ is not periodic, establishing the existence of periodic points is not immediate, but this part is less conceptual and more technical.
Using earlier results (cf. [Ruk22]), the Main Theorem can be reduced to the case of even $N$: for an odd $N$, there is a certain correspondence between outer billiards on the $N$-gon and on the $2N$-gon, allowing to conclude that the existence of aperiodic points in one of these systems implies the same for the other. Thus, it suffices to assume that $N$ is even. We also need to reduce the dynamics of $f_{\Pi}$ to that of a polygon exchange transformation, i.e., a piecewise isometry acting by translations. One obvious choice is the second iterate $f^{2}_{\Pi}$, however, a different choice allows to significantly simplify the computations. Namely, consider the half-turn $M$ about the center of $\Pi$, and form the composition $g:=M\circ f_{\Pi}$. It acts as a piecewise translation, see Fig. 3. Also, an $f_{\Pi}$-invariant bounded domain is invariant under $M$ and hence under $g$, and the dynamics of $g$ is not much different from that of $f_{\Pi}$, since $f_{\Pi}$ commutes with $M$. For these reasons, if $g$ is not periodic, then neither is a bounded restriction of $f_{\Pi}$. To establish non-periodicity of $g$, we use a two-dimensional dynamic invariant based on the Hadwiger–Glur valuation [HG51]. Explicit computation of this invariant and verification that the result is nonzero are not hard.
@@
Even though the two approaches are, by the end of the day, rely on similar ideas (non-vanishing of dynamic invariants), their implementations look different, and the invariants employed are also different. Any one of the two approaches is enough to prove the main theorem, however, we describe both. This is done on purpose: one or the other approch may work better for other problems.
@@ 1.3. Organization of the paper
Section 2 discusses general concepts associated with scissors congruences, including their dynamical aspects. As a special case, we consider polygon exchange transformations, which act as piecewise translations. Dynamic invariants of polytope exchange transformations of a given polygon are certain homomorphisms of the group of such transformations to the additive group of a real vector space. It follows that these dynamic invariants are conjugacy invariants under orientation preserving piecewise Euclidean conjugacies, and, for this reason, they are also called dynamic invariants. Dynamic Hadwiger invariants can be obtained as special cases of more general dynamic invariants; they are based on (translation) scissors congruence invariants of Hadwiger and Glur [HG51] initially discovered as a means of solving translation scissors congruence problem for planar polygons.
Assuming that $N$ is even, Section 3 deals with the outer billiard map $f_{\Pi}$ on a regular $N$-gon $\Pi$. There is a bounded domain $X$ invariant under $f_{\Pi}$; see Section 3.1. As explained above, the map $g_{X}:=M\circ f_{\Pi}|_{X}$ is a polygon exchange transformation of a bounded domain. Here $M$ is the half-turn about the center of $\Pi$. Computing the dynamic Hadwiger invariant of $g_{X}$ (Theorem 3.6) and establishing that it is nonzero (Corollary 3.7) yields that $f_{\Pi}$ is not periodic on $X$ (Theorem 3.8). We still need to upgrade this result to existence of aperiodic points, which is done in Section 6.
@@ 1.4. Acknowledgements
Valery Zgursky was a member of the team at an earlier stage of the project and co-authored a previous version of the paper; with gratitude, we acknowledge his contribution to the overall understanding of the problem and intuition through a series of computer experiments. We are grateful to Victor Kleptsyn for fruitful discussions and for communicating a proof of Proposition 2.33, to Alexandra Skripchenko for consultations on the topic of interval exchange transformations.
@@ 2.1. Basic terminology
Given a scissors isomorphism $f$ of $\Delta$, define a boundary point of $f$ as a point $z\in\overline{\Delta}$, for which there is $n\in\mathbb{Z}$ such that $z\notin\mathrm{dom}(f^{n})$, that is, some positive or negative iterate of $f$ is not defined at $z$ (which includes the case $z\in\partial\Delta$). Clearly, boundary points of $f$ are precisely boundary points of all forward or backward pieces of $f$. If $f^{n}(z)=z$ for some $n\neq 0$ (which means, in particular, that $z\in\mathrm{dom}(f^{n})$), then the point $z$ is called periodic under $f$. The smallest positive integer $n$ with this property is called the period of $z$. It is immediate from the definition that periodic points are not boundary points. Aperiodic points are defined as points of $\Delta$ that are neither periodic points nor boundary points.
@@ 2.2. Fatou and Julia sets
The Fatou set is cut by the set of boundary points into open domains which are either periodic or wandering (an open domain $\Omega\subset F(f)$ is said to be wandering if $f^{n}(\Omega)\cap\Omega=\varnothing$ unless $n=0$). The latter case is of course impossible if $\Delta$ has finite area.
@@
Clearly, if $f$ is periodic, then there are no essential open sets, hence $J$ is empty. On the other hand, if $f$ is non-periodic, then there is either an aperiodic point, which necessarily lies in $J$, or a sequence of periodic points whose periods tend to infinity, and any partial limit of this sequence, i.e., the limit of any converging subsequence, is in $J$. ∎
@@
Recall that $f$ preserves the Euclidean area, in particular, the Poincaré recurrence theorem is applicable to $f$. It says that all points of $\Delta$ are recurrent, except possibly for a set of zero area.
@@
Suppose that $f$ is periodic of period $n_{*}$, then $\Gamma_{+}=\Gamma_{n_{*}}$. Conversely, assume that $\Gamma_{+}=\Gamma_{n_{*}}$ for some $n_{*}>0$. The complement of $\Gamma_{+}$ is forward invariant under $f$, hence all complementary components are periodic. It follows that the complement of $\Gamma_{+}$ is also backward invariant under $f$, that is, $\Gamma_{-}=\Gamma_{+}=\Gamma$, as claimed. ∎
@@ 2.3. Polygon exchange transformations
Let $\mathsf{T}$ be the group of all translations of the plane; $\mathsf{T}$-scissors automorphisms are also known as translation scissors congruences or as polygon exchange transformations. The notion of a polygon exchange transformation admits a straightforward generalization to higher dimensions, where one considers polytope exchange transformations. Dimension one polytope exchange transformations are called interval exchange transformations. An interval exchange transformation can be defined on a straight line interval or a finite union of intervals.
Given a polytope, i.e., a bounded polyhedron, $\Delta$ of dimension $d$, we write $\mathrm{PET}(\Delta)$ for the group of polytope exchange transformations of $\Delta$, that is, the group of all automorphisms of $\Delta$ in the $\mathsf{T}$-scissors congruence groupoid. More precisely, elements of $\mathrm{PET}(\Delta)$ are classes of polytope exchange transformations with respect to the equivalence relation identifying two maps if they coincide except on a finite union of affine hyperplanes. Periodic polytope exchange transformations yield, in particular, periodic elements of the group $\mathrm{PET}(\Delta)$. If $d=1$, then one writes $\mathrm{IET}(\Delta)$ for the group of all interval exchange transformations of $\Delta$. Abbreviations IET and PET stand, of course, for “interval exchange transformations” and “polytope exchange transformations”, respectively. We concentrate on dimensions two or one in this paper even though some notions and results make sense for arbitrary dimensions.
@@
Let $I$ be an interval and $I_{1}$, $I_{2}\subset I$ be disjoint (except possibly a shared endpoint) subintervals of equal length. Then there is an interval exchange transformation $s$ that takes $I_{1}$ onto $I_{2}$, takes $I_{2}$ onto $I_{1}$, and leaves all other points of $I$ fixed; of course, $s$ may remain undefined on the common endpoint of $I_{1}$ and $I_{2}$. Call $s$ an interval swap map.
@@
A point of $\Delta$ can lie in the impressions of only finitely many nests, in fact, no more than $m_{*}$ nests, where $m_{*}$ is the same as in Lemma 2.13.
@@
Lemma 2.13 also motivates the notion of an elementary sector.
@@
    any sector with property (1) includes $\Lambda_{z}$ or is disjoint from $\Lambda_{z}$.
Thus, elementary sectors are minimal possible sectors formed by bi-pieces.
@@
Here, $m_{*}$ is again the same as in Lemma 2.13. For example, if $z$ is not a boundary point of $f$, then it lies in the impression of only one nest (namely, the level $n$ component of this nest is recovered as the level $n$ bi-piece containing $z$), and the only elementary sector at $z$ is the entire plane.
2.4. Sah–Arnoux–Fathi invariants
We proceed by reviewing an invariant of an interval exchange transformation called the Sah–Arnoux–Fathi invariant (abbreviated as SAF invariant) [Arn81, Sah81]. Consider an interval exchange transformation $g$ defined on an interval $I$. Then the SAF invariant $\mathrm{Inv}(g)$ is an element of $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ defined as follows. If subintervals $I_{1}$, $\dots$, $I_{n}$ of lengths $\lambda_{1}$, $\dots$, $\lambda_{n}$, respectively, form a partition of $I$, and the restriction of $g$ to each $I_{i}$ is a translation by $\tau_{i}\in\mathbb{R}$, then
$\mathrm{Inv}(g)=\lambda_{1}\otimes\tau_{1}+\lambda_{2}\otimes\tau_{2}+\dots+\lambda_{n}\otimes\tau_{n}.$
@@ 2.4. Sah–Arnoux–Fathi invariants
$\mathrm{Inv}_{\beta}(g):=\beta(\lambda_{1},\tau_{1})+\beta(\lambda_{2},\tau_{2})+\dots+\beta(\lambda_{n},\tau_{n}).$
@@
Next, we want to generalize the SAF invariant to higher dimensional polytope exchange transformations. Specifically, define a dynamic invariant as a group homomorphism from $\mathrm{PET}(P)$ to $\mathbb{R}$. In what follows, new dynamic invariants will be introduced in 2D (they also make sense in dimensions $>2$).
@@ 2.5. The ideal Julia set
A nest $(Q_{n})$ is said to be essential if $Q_{n}$ is an essential set (see Definition 2.6), for every $n\geqslant 0$. Say that $(Q_{n})$ is periodic if $Q_{n+1}=Q_{n}$ for all sufficiently large $n$; otherwise $(Q_{n})$ is said to be non-periodic.
@@
All results about the ideal Julia set can be rephrased in terms of symbolic dynamics. Consider the alphabet $\mathcal{A}$ whose symbols are in one-to-one correspondence with components of $\mathrm{dom}(f)$; one can formally set $\mathcal{A}:=\pi_{0}(\mathrm{dom}(f))$. Every nest $(Q_{n})$ gives rise to a bi-infinite sequence in the alphabet $\mathcal{A}$, namely, the $n$-th symbol is the component of $\mathrm{dom}(f)$ containing $f^{n-1}(Q_{n})$ if $n>0$ and the component of $\mathrm{dom}(f)$ containing $f^{n-1}(Q_{-n+1})$ if $n\leqslant 0$. This correspondence embeds the dynamical system $(\mathcal{J},h_{f})$ into $(\mathcal{A}^{\mathbb{Z}},\sigma)$, where $\sigma$ is the shift map.
@@ 2.6. Valuations
Let $\mathcal{P}$ be the set of all compact convex polygons in the plane, including $\varnothing$. A valuation on convex polygons with values in $\mathbb{R}$ is by definition a function $\mu:\mathcal{P}\to\mathbb{R}$ such that $\mu(\varnothing)=0$ and
@@
Fix an area form on $\mathbb{R}^{2}$. Then one can assign the area $\mathrm{area}(P)$ to any polytope $P$. Area is a simple valuation invariant under the full group of isometries of $\mathbb{R}^{2}$. Now let $\xi:\mathbb{R}\to\mathbb{R}$ be any $\mathbb{Q}$-linear function. Then $\xi\circ\mathrm{area}$ is also a simple isometry invariant valuation; this construction gives uncoutably many such valuations, linearly independent over $\mathbb{R}$. Perimeter (suitably extended to degenerate polygons) is also an isometry invariant valuation, however, it is not simple.
@@
Let $P$ be a bounded polygon. By a dynamic invariant, we mean a group homomorphism $\Psi$ from $\mathrm{PET}(P)$ to the additive group of some $\mathbb{Q}$-vector space. For any two transformations $f$, $g\in\mathrm{PET}(P)$, we therefore have
@@
It follows that $\Psi(f^{n})=n\Psi(f)$ for every $n\in\mathbb{Z}$. Therefore, if $f$ is periodic, then any dynamic invariant must vanish on $f$. Dynamic invariants are, indeed, congugacy invariants ($\Psi(g\circ f\circ g^{-1})=\Psi(f)$). Below, we give a general construction of multidimensional dynamic invariants.
@@
$\mathrm{Inv}_{\mu}(f)=\sum_{i=1}^{n}\mu(P_{i})\otimes\vec{v}_{i}\in\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{2}$
@@
In particular, any $\mathbb{Q}$-bilinear functional $\beta:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{C}$ defines a dynamic invariant $\mathrm{Inv}_{\mu,\beta}$ with values in $\mathbb{C}$, which can be computed as the sum of $\beta(\mu(P_{i}),\vec{v}_{i})$ for $i=1$, $\dots$, $n$.
@@
The dynamic Hadwiger invariant lies in the kernel of the multiplication map $\mathfrak{m}:\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{2}\to\mathbb{R}^{2}$ given by $\mathfrak{m}(x\otimes\vec{v})=x\vec{v}$.
@@
Let $f$ be a polygon exchange transformation of a bounded polygon $P$ so that $\mathrm{dom}(f)$ consists of open polygons $P_{1}$, $\dots$, $P_{n}$, and $f|_{P_{i}}$ is the translation by vector $\vec{v}_{i}$. Observe that $\mathfrak{M}(f):=\mathfrak{m}(\mathrm{Inv}_{L}(f))$ can be computed as the sum of $\mu_{L}(P_{i})\vec{v}_{i}$. We want to show that $\mathfrak{M}(f)=0$.